OpenStudy (anonymous):

How do you factor 4x^3-13x^2+3x? I started by taking the GCF (x) out, and I got 4x^2-13x+3, but after that, I'm not sure what to do.

hero (hero):

If you factored out x then you have \[x(4x^2 - 13x + 3)\] Now find two numbers, \(m\) and \(n\) that add to get -13, yet multiply to get 12. \[m + n = -13 \\ m \times n = 12\]

OpenStudy (anonymous):

-12 and -1 I tried that, but then when I put them in factored for, I tried checking for the answer and it didn't work. (4x-1) (4x-12)

hero (hero):

Okay, I see that your factorization is not correct.

hero (hero):

Here's what you're supposed to do after finding the two numbers \(x(4x^2 -13x + 3)=x(4x^2 -12x - 1x + 3)\) \(=x(4x(x - 3) -1(x - 3))\) \(=x((x - 3)(4x - 1))\) \(=x(x - 3)(4x - 1)\)

OpenStudy (anonymous):

THANK YOU! I wasn't taught that method, but now that you've shown me, THANKS! If this were any other problem where I didn't have a GCF, could I still keep the -12x and -1x separate?

hero (hero):

Yes. The method is called factor by grouping. If the quadratic is factorable, then you can use that method. However, you have to know what you're doing in order to use it.

OpenStudy (anonymous):

Yup, I understand, thank you!

OpenStudy (anonymous):

Quick question, sorry to come back to this problem again, but the -1x+3... when you factored the -1 out, are you allowed to change the sign from a + to a - inside the parentheses of the x+3?

hero (hero):

Basically.... Suppose you had \[-1(x - 3)\] And you were asked to use the distributive property to expand the expression...what would be the result?

hero (hero):

If you were to expand the expression using the distributive property, you would see that you arrive back at the original expression which is \(-1x + 3\) because \(-1(x - 3) = -1(x) -1(-3)= -1x + 3\)